OpenStudy (anonymous):

What would the simple interest rate be if you earned $136 on a $1,700 investment in one year.?

OpenStudy (slaaibak):

Solve for i: 1700i = 136

OpenStudy (slaaibak):

The usual formula is: A = P(1 + ni) where A is the accumulation, P the initial amount, n the time in years and i the interest rate. We can show it works using this formula: 1836 = 1700(1 + i) 1836 = 1700 + 1700i 136 = 1700i which is the same as above

OpenStudy (anonymous):

I=PRT is the simple interest rate formula I= Interest earned P=Premium R=Rate T=Time Plug it in and you get your answer. The above"A = P(1 + ni)" also give you the same answer.